OpenStudy (gudetamas):

if l||m, m<4 = 12x + 5, and m<5 = 8x+17, then what is m<2?

OpenStudy (gudetamas):

i know you start by 12x+5 = 8x +17 and get 20x+22

jimthompson5910 (jim_thompson5910):

can you draw the picture of how angles 2,4,5 are set up?

OpenStudy (gudetamas):



jimthompson5910 (jim_thompson5910):

thanks

jimthompson5910 (jim_thompson5910):

you are correct in saying that m < 4 = m < 5 which leads to 12x+5 = 8x +17

jimthompson5910 (jim_thompson5910):

then 12x+5 = 8x +17 turns into 4x = 12 I think that's what you meant to say? I subtracted 8x from both sides. I also subtracted 5 from both sides

jimthompson5910 (jim_thompson5910):

if 4x = 12, then x = ???

OpenStudy (gudetamas):

3?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

then use x = 3 to compute the value of m < 4

OpenStudy (gudetamas):

so 41?

jimthompson5910 (jim_thompson5910):

yep and angle 2 is equal to that (since angles 2 and 4 are vertical angles, they are congruent)